# DynaSeg: A Deep Dynamic Fusion Method for Unsupervised Image Segmentation Incorporating Feature Similarity and Spatial Continuity

### v1

@@ Abstract
Our work tackles the fundamental challenge of image segmentation in computer vision, which is crucial for diverse applications. While supervised methods demonstrate proficiency, their reliance on extensive pixel-level annotations limits scalability. In response to this challenge, we present an enhanced unsupervised Convolutional Neural Network (CNN)-based algorithm called DynaSeg. Unlike traditional approaches that rely on a fixed weight factor to balance feature similarity and spatial continuity, requiring manual adjustments, our novel, dynamic weighting scheme automates parameter tuning, adapting flexibly to image details. We also introduce the novel concept of a Silhouette Score Phase that addresses the challenge of dynamic clustering during iterations. Additionally, our methodology integrates both CNN-based and pre-trained ResNet feature extraction, offering a comprehensive and adaptable approach. We achieve state-of-the-art results on diverse datasets, with a notable 12.2% and 14.12% mIOU improvement compared to the current benchmarks on COCO-All and COCO-Stuff, respectively. The proposed approach unlocks the potential for unsupervised image segmentation and addresses scalability concerns in real-world scenarios by obviating the need for meticulous parameter tuning.
Index Terms:
Unsupervised Learning, Image Segmentation.
I Introduction
@@
In recent years, computer vision has witnessed remarkable strides, driven by the availability of large-scale image datasets as well as the development of advanced machine learning algorithms. Within various real-time applications[1][2][3][4], image segmentation stands out as a pivotal computer vision task, playing a crucial role in interpreting visual content at a granular level. Unlike the straightforward task of image classification, which allocates a category label to the entire image, segementation involves assigning category labels to individual pixels, effectively outlining the image into meaningful segments or regions. This nuanced approach not only enhances the interpretation of complex scenes, but also contributes to a deeper understanding of visual content.
@@
The segmentation task can be approached from different perspectives, leading to three main types of image segmentation. Semantic segmentation[5] identifies uncountable and shapeless regions, often referred to as ’stuff,’ such as grass, sky, or road, based on similar textures or materials. Instance segmentation[6] focuses on pinpointing and segmenting individual instances of countable objects, such as people, animals, and tools, treating them as distinct ’things.’ Lastly, panoptic segmentation[7] unifies the two distinct concepts used to segment images, assigning a semantic label to each pixel in an image and giving the pixel a unique identifier if it is an object instance. The semantic nuances that differentiate these tasks have led to the creation of methods that employ specialized architectures. Although these task-specific architectures have significantly advanced their segmentation objectives, they need more flexibility to generalize across different segmentation tasks. Variability in the object’s appearance, complex interaction with the object, variations in the object’s scale, and noisy or ambiguous edges complicate the process further.
Classical pioneering segmentation techniques such as active contour models (ACM) [8], k-means [9], and graph-based segmentation (GS) [10][11] impose global and local data and geometry constraints on the masks. As a result, these techniques are sensitive to initialization and require heuristics such as point resampling, making them unsuitable for modern applications.
@@
The state of the art in segmentation is dominated by deep learning. The Mask R-CNN framework [12] has provided advancement in semantic and instance image segmentation. However, a drawback lies in the substantial requirement for hand-labeled data, limiting the widespread applicability across diverse domains. This challenge becomes particularly pronounced in the context of pixel-wise classification, where the cost of annotation per image is prohibitively expensive.
Unsupervised image segmentation is a possible solution to automatically segment an image into semantically similar regions where the system can find objects, precise boundaries and materials that even the annotation system may not unveil properly. The task has been studied as a clustering problem in the recent literature, with promising results.
@@
Differentiable feature clustering, a state-of-the-art CNN-based algorithm proposed by [13], simultaneously optimizes pixel labels and feature representations through a combination of feature similarity and spatial continuity constraints [13]. Feature similarity corresponds to the constraint that pixels in the same cluster should be similar to each other. Spatial continuity refers to the constraint that pixels in the same cluster should be next to each other (continuous). However, to achieve the desired segmentation result, [13] applies a manual parameter tuning to find the optimal balancing weight $\mu$, which fails to achieve a good balance between the two constraints, as mentioned above, depending on the degree of detail in the image and the dataset.
@@
  • Dynamically Weighted Loss: Introducing a flexible and adaptive loss function that dynamically adjusts the balancing weight $\mu$ during training.
  • ResNet with FPN Integration: Leveraging the power of ResNet with Feature Pyramid Network (FPN) for enhanced feature extraction in the unsupervised segmentation task.
  • Silhouette Score Phase: Incorporating a silhouette score-based phase to guide the optimization process, ensuring improved cluster quality.
@@
II Related Work
Machine learning techniques have achieved remarkable success in various vision applications, notably in classification [15] and object detection [16]. Image segmentation in particular has received a lot of research attention.
II-A Supervised Segmentation
Recent advances in convolutional neural networks (CNN) and the availability of large annotated datasets have outperformed traditional algorithms [10, 8], which face challenges such as susceptibility to noise, difficulty handling intricate boundaries, and sensitivity to data quality [17, 18, 19]. Techniques like U-Net [20], SegNet [21], and DeepLab [22] illustrate the effectiveness of CNNs in achieving precise segmentation.
L. Jonathan et al. in [23], inspired by the VGG network [24], introduced FCN, a fully convolutional network with multi-resolution layers for segmentation. Despite these innovations, the FCN result remains blurry, smooth, and unsatisfactory. Afterward, O. Ronneberger in [20] proposed a U-net autoencoder with fewer annotated images and a reasonable training time, addressing the scarcity of training images in biomedical applications.
In [12], Kaiming et al. present Mask R-CNN, an extension of Faster R-CNN [12], incorporating a branch for predicting an object mask alongside the existing branch for bounding box recognition. Following the success of Mask R-CNN, several deep neural architectures have been proposed recently, such as transformers [25] and capsule networks [26], showing promising results. However, their reliance on hand-labelled data escalates annotation costs and introduces challenges related to missing labels, small datasets, and low resolution [27, 28, 29].
While supervised segmentation provides excellent results om benchmark datasets, real-life application becomes costly due to its heavy dependence on annotated data. This inherent limitation has resulted in the rise of recent unsupervised segmentation models, including our proposed model, which operates effectively in the absence of explicit annotations, addressing the need for scalable and cost-efficient segmentation solutions.
@@ II-B Unsupervised Segmentation
Unsupervised segmentation methods have recently emerged as a cost-effective alternative, aiming to discover objects, boundaries, and materials without explicit ground truth. We divide our discussion into two subsections discussing clustering and CNN-based approaches.
II-B1 Unsupervised Clustering for Segmentation
At the forefront of unsupervised segmentation methodologies, Invariant Information Clustering (IIC) [30], and its versions [31] operate at the edge level, training a neural network to form semantic clusters without the need for postprocessing. However, its reliance on enforcing a uniform distribution over clusters makes it particularly effective with well-balanced datasets [28, 29]. Yet, the efficacy of domain knowledge-based approaches, including IIC [30], is contingent upon data balance and can be influenced by the specific data augmentation strategy employed [32, 33, 34].
To address the challenges associated with IIC and similar techniques, unsupervised segmentation, traditionally framed as a clustering task, leverages methods like K-means [9] and graph-based segmentation (GS) [10]. These approaches utilize features such as color and pixel information to discern meaningful structures within the data. However, K-means struggles with continuous areas due to its inherent sharp boundary assumptions, while GS strikes a balance between preserving details in low-variability regions and ignoring high-variability ones across image regions.
Inspired by the principles of unsupervised segmentation methods, our model adopts a similar clustering strategy. By optimizing clustering loss, we aim to enhance the feature space, accounting for high-level visual similarity. Importantly, our model eliminates the need for a postprocessing phase, streamlining the segmentation process and offering a more efficient and effective solution.
II-B2 CNN-based Unsupervised Approaches
Recent advancements in deep learning and convolutional neural architectures enable effective clustering in well-embedded feature spaces. CNN-based unsupervised approaches can be categorized into two groups: discriminative models and generative models.
Generative Models
Generative Adversarial Networks (GANs) map latent codes to images, generating foreground and background separately. Melas-Kyriazi [35] proposes a GAN-based method for unsupervised foreground-background separation. Leveraging pre-trained GANs like BigBiGAN [36], this approach extracts salient object segmentation from the latent space. However, it relies on class supervision. Labels4Free [37] augments Style-GAN2 [38], introducing a segmentation branch for unsupervised foreground object segmentation. These generative models often focus on binary clustering, distinguishing a specific foreground object.
Discriminative Models
In contrast to generative models, discriminative models focus on multi-classification. Wnet [39] exemplifies this approach by combining two U-net [20] structures into an auto-encoder, followed by a postprocessing phase for segmentation refinement. While effective, this method is computationally expensive and requires extensive hyperparameter tuning.
Another discriminative model, Deep Image Clustering (DIC) [40], takes a cost-effective route by breaking down the segmentation problem into two steps: a feature transformation sub-network and a trainable deep clustering sub-network. However, it relies on superpixels, introducing predetermined boundaries.
Infomax [41] presents an alternative where an input image is partitioned into superpixels. The Region-Wise-Embedding (RWE) extracts a feature embedding for each superpixel region, and Mutual-Information-Maximization is employed with adversarial training. Despite its efficiency, challenges persist in recognizing classes due to inherent similarities.
Moving beyond traditional methods, Pixel-wise Clustering Information Extraction (PICIE) [42] introduces a Siamese perspective to unsupervised segmentation, emphasizing equivariance learning. This ensures consistent clustering assignments across different image views and mitigates issues related to hyperparameter sensitivity and dataset imbalance.
Inspired by PICIE, DenseSiam [43] adopts a focus on learning dense representations through contrastive learning and clustering-based techniques, incorporating the class-balanced cross-entropy loss. However, these methods introduce additional computational operations, potentially struggling with unknown transformations.
In contrast to approaches relying on intermediate representations followed by postprocessing [39, 40, 41] or being limited to known transformations [43, 42], our alternative strategy, inspired by Kim et al. [13], jointly optimizes pixel labels and features. This process, utilizing the backpropagation of the feature similarity loss and a spatial continuity loss, efficiently generates compact clusters. We emphasize dynamic optimization of the clustering loss to enhance the feature space, addressing challenges posed by superpixels in [44] and the data sensitivity associated with weight coefficients in [13].
III Methodology
III-A Model Architecture
@@
Our unsupervised image segmentation model, as illustrated in Figure (1), is designed to capture complex patterns and structures within images effectively. Building upon the architecture introduced in our preliminary work [14], we present further enhancements tailored to adapt to varying image content. The architecture comprises three main components: a Feature Extractor Network, a Dynamic Weighting Scheme, and a Clustering Mechanism.
@@
The model, DynaSeg, is shown in Figure (1), where the Feature Extractor Network is employed to produce a $p$-dimensional feature map $r$. A final linear classifier layer classifies the features of each pixel into $q$ classes. A batch normalization function is applied to the response map $r$ to obtain a normalized map $r^{\prime}$. Lastly, the argmax function is utilized to select the dimension that has the maximum value in $r^{\prime}$. Each pixel is assigned the corresponding cluster label $c_{n}$, identical to allocating each pixel to the closest point among the $q^{\prime}$ representative points.
@@
During backward propagation, the loss $L$ (defined in the next section) is calculated. Subsequently, the convolutional filters’ parameters and the classifier’s parameters are updated using stochastic gradient descent. This forward-backward process is iterated $T$ times to obtain the final prediction of the cluster labels $c_{n}$. The segmentation problem is handled in an unsupervised manner without knowing the exact number of clusters, making it flexible according to the content of the image. Initially, a large number $q$ is allocated to the initial cluster labels $q^{\prime}$. Subsequently, similar or spatially related pixels are iteratively integrated to update the number of clusters $q^{\prime}$.
The Feature Extractor Network is the initial processing stage responsible for extracting high-level features from input images. In our extended model, we explore two alternatives for feature extraction: a CNN-based Network comprising $M$ convolutional components and a pre-trained ResNet [45] with a decoder. This exploration allows us to compare and leverage the strengths of each approach for improved image segmentation.
Simultaneously, the Dynamic Weighting Scheme dynamically adjusts the weighting parameter ($\mu$) during iterations, introducing two variants: Feature Similarity Focus (FSF) and Spatial Continuity Focus (SCF). This adaptive scheme enables the model to shift its focus between feature similarity and spatial continuity, reducing sensitivity to parameter tuning and enhancing overall segmentation performance.
The Clustering Mechanism assigns cluster labels to pixels based on the features extracted by the network. The clustering is performed iteratively, with the number of clusters dynamically updated based on feature similarity and spatial continuity. The initial (maximum) number of cluster labels $q$ is set to a large value, and through iterative updates, similar or spatially close pixels are integrated, reducing the number of unique cluster labels $q^{\prime}$. To prevent under-segmentation, an intra-axis normalization process is introduced, giving preference to a larger $q^{\prime}$.
As a novel addition to our methodology, we introduce the Silhouette Score Phase, calculated only once at the first iteration of the raw image. The optimal number of clusters determined by the silhouette score serves as a threshold, preventing convergence to a single cluster and ensuring meaningful segmentation.
@@
III-B Feature Extractor Network

### v3

@@ Abstract
Our work tackles the fundamental challenge of image segmentation in computer vision, which is crucial for diverse applications. While supervised methods demonstrate proficiency, their reliance on extensive pixel-level annotations limits scalability. We introduce DynaSeg, an innovative unsupervised image segmentation approach that overcomes the challenge of balancing feature similarity and spatial continuity without relying on extensive hyperparameter tuning. Unlike traditional methods, DynaSeg employs a dynamic weighting scheme that automates parameter tuning, adapts flexibly to image characteristics, and facilitates easy integration with other segmentation networks. By incorporating a Silhouette Score Phase, DynaSeg prevents undersegmentation failures where the number of predicted clusters might converge to one. DynaSeg uses CNN-based and pre-trained ResNet feature extraction, making it computationally efficient and more straightforward than other complex models. Experimental results showcase state-of-the-art performance, achieving a 12.2% and 14.12% mIOU improvement over current unsupervised segmentation approaches on COCO-All and COCO-Stuff datasets, respectively. We provide qualitative and quantitative results on five benchmark datasets, demonstrating the efficacy of the proposed approach. Code available at https://github.com/RyersonMultimediaLab/DynaSeg
1 Introduction
@@
In recent years, computer vision has taken remarkable strides, driven by the availability of large-scale image datasets as well as the development of advanced machine learning algorithms. Within various real-time applications [1][2][3][4], image segmentation stands out as a pivotal computer vision task, playing a crucial role in interpreting visual content at a granular level. Unlike the straightforward task of image classification, which allocates a category label to the entire image [5], segmentation involves assigning category labels to individual pixels, effectively outlining the image into meaningful segments or regions. This nuanced approach not only enhances the interpretation of complex scenes but also contributes to a deeper understanding of visual content.
@@
The segmentation task can be approached from different perspectives, leading to three main types of image segmentation. Semantic segmentation [6] identifies uncountable and shapeless regions, often referred to as ’stuff,’ such as grass, sky, or road, based on similar textures or materials. Instance segmentation [7] focuses on pinpointing and segmenting individual instances of countable objects, such as people, animals, and tools, treating them as distinct ’things.’ Lastly, panoptic segmentation [8] unifies the two distinct concepts used to segment images, assigning a semantic label to each pixel in an image and giving the pixel a unique identifier if it is an object instance. The semantic nuances that differentiate these tasks have led to the creation of methods that employ specialized architectures. However, despite the advancements in segmentation methodologies, significant challenges persist. Variability in object appearance, complex interactions between objects, variations in object scale, and noisy or ambiguous edges all pose obstacles to accurate segmentation. Objects can look different due to lighting, colour, and texture changes, making consistent identification difficult. Complex interactions, such as overlapping people in a crowd or intertwined tree branches and power lines, further complicate segmentation. Additionally, objects appear in various sizes depending on their distance from the camera, requiring algorithms to handle different scales effectively. Noisy or ambiguous edges, caused by low resolution or motion blur, add another layer of difficulty. Developing robust and adaptable segmentation methods that can address these complexities is crucial for improving the accuracy and reliability of image segmentation in real-world applications. Therefore, while task-specific architectures have significantly advanced segmentation objectives, there is a pressing need for methods that offer more flexibility to generalize across different segmentation tasks, ultimately enhancing the performance of segmentation algorithms in challenging real-world scenarios.
@@
Classical pioneering segmentation techniques such as active contour models (ACM) [9], k-means [10], and graph-based segmentation (GS) [11][12] impose global and local data and geometry constraints on the masks. As a result, these techniques are sensitive to initialization and require heuristics such as point resampling, making them unsuitable for modern applications. Currently, the state of the art in segmentation is dominated by deep learning. The Mask R-CNN framework [13] has provided advancement in semantic and instance image segmentation. However, a major drawback is a substantial requirement for hand-labelled data, limiting the widespread applicability across diverse domains. This challenge becomes particularly pronounced in the context of pixel-wise classification, where the cost of annotation per image is prohibitively expensive.
@@
Unsupervised image segmentation is a possible solution to automatically segment an image into semantically similar regions where the system can find objects, precise boundaries and materials that even the annotation system may not unveil properly. The task has been studied as a clustering problem in the recent literature [14][15][16], with promising results. Differentiable feature clustering, a state-of-the-art CNN-based algorithm proposed by [17], simultaneously optimizes pixel labels and feature representations through a combination of feature similarity and spatial continuity constraints [17]. Feature similarity corresponds to the constraint that pixels in the same cluster should be similar to each other. Spatial continuity refers to the constraint that pixels in the same cluster should be next to each other (continuous). However, to achieve the desired segmentation result, [17] applies a manual parameter tuning to find the optimal balancing weight $\mu$, which fails to achieve a good balance between the two constraints, as mentioned above, depending on the degree of detail in the image and the dataset.
@@
  • Dynamically Weighted Loss: Introducing a flexible and adaptive loss function that dynamically adjusts the balancing weight $\mu$ during training, ensuring optimal balance between feature similarity and spatial continuity. This makes the model adaptable to different datasets without requiring manual changes to parameters, such as the balancing weight, and facilitates easy implementation of the proposed method on other segmentation networks.
  • Silhouette Score Phase: Incorporating a silhouette score-based phase to guide the optimization process, ensuring improved cluster quality. Unlike traditional methods that rely on fixed thresholds to determine segmentation termination and may converge to a single cluster, failing to capture the diversity and complexity of real-world images, our approach dynamically assesses cluster quality using silhouette scores. This prevents under-segmentation failures, where the segmentation process prematurely stops, leading to inaccurate results. By dynamically evaluating cluster quality, our method ensures that the segmentation process continues until meaningful clusters are formed, without the need for predefined thresholds. This not only improves segmentation accuracy but also reduces the burden of fine-tuning additional hyperparameters.
  • ResNet with FPN Integration: Leveraging the power of ResNet combined with the Feature Pyramid Network (FPN) for enhanced feature extraction in the unsupervised segmentation task. The FPN enhances the representation map, making it semantically strong and improving the overall segmentation performance. This integration is particularly beneficial for segmenting objects at different scales and resolutions, as FPN enables multi-scale processing and feature integration across different levels of abstraction. Additionally, by incorporating ResNet, we capitalize on its inherent strengths, similar to those of CNNs, while leveraging residual information to mitigate issues like vanishing gradients. This combination amplifies the network’s capacity to capture intricate details and semantic information critical for accurate segmentation.
@@
2 Related Work
Recent advancements in machine learning have propelled various vision applications, notably in classification [19] and object detection [20]. Image segmentation, a crucial task in computer vision, has seen significant progress, primarily driven by supervised learning techniques. However, the reliance on annotated data poses challenges in scalability and cost. This section reviews existing methods and situates our approach within the broader landscape, emphasizing its strengths and addressing the limitations of previous techniques.
2.1 Supervised Segmentation
The advent of convolutional neural networks (CNNs) and large annotated datasets has revolutionized image segmentation, surpassing traditional algorithms such as active contour models (ACM) [9], k-means [10], and graph-based segmentation (GS) [11]. Traditional methods often struggle with noise susceptibility, intricate boundary handling, and sensitivity to data quality [21, 22, 23]. In contrast, CNN-based techniques like U-Net [24], SegNet [25], and DeepLab [26] offer precise segmentation capabilities.
Among the pioneering approaches, the Fully Convolutional Network (FCN) [27] inspired by the VGG network [28] utilized multi-resolution layers for segmentation. However, FCN results often appeared blurry and smooth, needing more precision. Addressing this, Ronneberger proposed U-Net [24], catering to fewer annotated images with reasonable training times, notably in biomedical applications. Further advancements were made with Mask R-CNN [13], extending Faster R-CNN [29] by predicting object masks alongside bounding boxes, significantly improving instance segmentation performance. Following the success of Mask R-CNN, several novel deep neural architectures have been proposed, including transformers [30] and capsule networks [31], both showing promising results. Despite their effectiveness, these models rely heavily on large annotated datasets, escalating annotation costs and introducing challenges related to missing labels and data quality variations [32, 33, 34].
While supervised segmentation methods achieve excellent results on benchmark datasets, their application in real-world scenarios is costly due to the dependence on extensive annotated data. This limitation has catalyzed the development of unsupervised segmentation models, including our proposed method. Our approach provides a scalable and cost-efficient solution, effectively addressing the need for robust segmentation without the necessity for explicit annotations.
@@ 2.2 Unsupervised Segmentation
Recent advancements in unsupervised segmentation methods offer a cost-effective alternative, aiming to discover objects, boundaries, and materials without explicit ground truth. Among these approaches, clustering-based and CNN-based methods have gained prominence, each offering unique strategies to tackle the challenges of unsupervised segmentation.
2.2.1 Clustering-Based Methods
Foundational unsupervised segmentation methods like K-means [10], Mean Shift [35], and graph-based segmentation (GS) [11] form clusters based on pixel feature similarity but often lack spatial coherence. These methods treat each pixel independently, resulting in fragmented segmentations with disconnected regions, especially images with complex structures and textures. In contrast, our model incorporates both feature similarity and spatial continuity to ensure that similar and contiguous pixels belong to the same cluster. Unlike static clustering methods [10][35] that require the number of clusters to be specified a priori, DynaSeg adjusts the number of clusters during training, enhancing the model’s flexibility and robustness.
One significant clustering-based approach is Deep Image Clustering (DIC) [15], which breaks down the segmentation problem into feature transformation and trainable deep clustering sub-networks. However, it relies on superpixels, introducing predetermined boundaries. Similarly, Invariant Information Clustering (IIC) [16], and its versions [36] uses mutual-information-maximization for clustering. However, its reliance on enforcing a uniform distribution over clusters makes it particularly effective with well-balanced datasets [33, 34]. Yet, the efficacy of IIC [16] is contingent upon data balance and can be influenced by the specific data augmentation strategy employed [37, 38, 39]. Unlike the methods [15][16] that separate feature extraction and clustering into distinct steps, DynaSeg employs Joint Optimization where the backpropagation of clustering losses directly influences the CNN. This integration ensures a cohesive segmentation process that is continuously refined, overcoming the limitations of predetermined boundaries and data imbalance.
2.2.2 Deep Learning-Based Methods
Deep learning-based methods have garnered significant attention in unsupervised image segmentation, showcasing remarkable capabilities across various approaches. For instance, Generative Adversarial Networks (GANs) [40] [41] [42] [43] have emerged as powerful tools for creating segmented images by effectively separating foreground and background. Specifically, the GAN-based approach proposed by Melas-Kyriazi [40] leverages pre-trained GANs like BigBiGAN [41] to extract salient object segmentation from the latent space. Despite their innovation, these methods often rely on some form of class supervision, which inherently limits their purely unsupervised nature. Labels4Free [42] extends Style-GAN2 [43] with a segmentation branch for unsupervised foreground object segmentation, typically focusing on binary clustering. However, these approaches tend to rely on discriminative losses, such as binary cross-entropy loss, for training the discriminator network, which may not directly optimize for segmentation quality, resulting in suboptimal outcomes. In contrast, our model adopts a hybrid approach for multi-class segmentation by integrating a clustering strategy with CNN-based feature extraction.
Beyond generative models, contrastive learning [44][45] has gained traction for its ability to learn representations by contrasting similar and dissimilar pairs of image patches. PICIE [44] introduces a Siamese perspective to unsupervised segmentation, emphasizing equivariance learning and consistent clustering assignments across different image views. DenseSiam [45] focuses on learning dense representations through contrastive learning and clustering-based techniques. While these methods effectively capture images’ inherent structures and patterns, contributing significantly to segmentation performance, they have notable limitations. Specifically, these approaches introduce additional computational operations, struggle with unknown transformations, and are sensitive to noisy or low-quality data, often resulting in biased representations or suboptimal clustering results. Our model addresses these limitations and offers a more robust solution. DynaSeg employs a spatial continuity loss that acts as a high-pass filter, enhancing segmentation accuracy by preserving high-frequency details and attenuating low-frequency components. This discourages smooth regions from being segmented into multiple clusters, promoting homogeneity within each cluster, making the segmentation more coherent and reducing noise within clusters.
Wnet [46] combines U-net [24] structures into auto-encoders for segmentation, followed by a post-processing phase for segmentation refinement. While effective, this method is computationally expensive and requires extensive hyperparameter tuning. Infomax [14] presents an alternative where an input image is partitioned into superpixels. The Region-Wise-Embedding (RWE) extracts a feature embedding for each superpixel region, and Mutual-Information-Maximization is employed with adversarial training. Despite its efficiency, it faces challenges like superpixel reliance and predetermined boundaries. In contrast, our proposed hybrid model offers a streamlined solution by eliminating the need for post-processing. Our model generates compact clusters directly from the feature space by dynamically optimizing the clustering loss and spatial continuity loss. This integrated approach enables our model to extract rich and informative features, facilitating accurate and robust segmentation, even in challenging scenarios. DynaSeg offers enhanced flexibility and adaptability by removing the reliance on predefined boundaries or superpixels [47], ensuring more natural and visually appealing segmentation outcomes.
Differentiable feature clustering, as proposed by [17], is a CNN-based algorithm that optimizes pixel labels and feature representations using a combination of feature similarity and spatial continuity constraints [17]. However, achieving the desired segmentation result with this approach often requires manual parameter tuning to find the optimal balancing weight ($\mu$) between feature similarity and spatial continuity. This manual tuning process can be cumbersome and may fail to balance the two constraints well, mainly when dealing with images of varying levels of detail and datasets with different characteristics. Moreover, relying solely on fixed weight factors for feature similarity and spatial continuity constraints can lead to the number of predicted clusters converging to one, making it challenging to adapt to different datasets and requiring extensive fine-tuning. Additionally, the lack of adaptability and the need for manual parameter adjustments make integrating this approach with other networks difficult. In contrast, our model introduces a Dynamic Weighting Scheme that dynamically adjusts the weighting parameter ($\mu$) during iterations. By incorporating a silhouette score-based phase to guide the optimization process, our model prevents under-segmentation failures and ensures robust performance across diverse datasets. This dynamic approach not only simplifies the integration of our model with other networks but also enhances its adaptability and scalability, making it well-suited for various segmentation tasks and datasets.
Additionally, recent advancements in CNN architectures, such as Vision Transformers (ViT) [48], have shown promise in various computer vision tasks, including image segmentation. However, ViTs come with their own set of challenges, such as high computational complexity, data inefficiency, training instability, and a propensity for overfitting, which can limit their accessibility and applicability in specific contexts. In contrast, our hybrid model offers a more efficient and scalable solution by dynamically optimizing the clustering loss and leveraging a CNN-based network or a pre-trained ResNet with FPN to extract high-level features from input images. This approach ensures robustness and adaptability across different datasets. Overall, our model provides a practical and effective solution for unsupervised image segmentation, addressing the limitations of existing methods and presenting promising opportunities for real-world applications.
3 Methodology
This section details the proposed DynaSeg method, including the architecture design, dynamic weighting scheme, silhouette score integration, and training procedure. Each component is meticulously described to ensure reproducibility and clarity.
3.1 Model Architecture
@@
Our unsupervised image segmentation model, DynaSeg, is designed to effectively capture complex patterns and structures within images. Building upon the architecture introduced in our preliminary work [18], we present further enhancements tailored to adapt to varying image content. The architecture comprises three main components: a Feature Extractor Network, a Dynamic Weighting Scheme, and a Clustering Mechanism.
@@
As shown in Figure (1), the Feature Extractor Network produces a $p$-dimensional feature map $r$, which is then classified into $q$ classes using a linear classifier layer. A batch normalization function is applied to obtain a normalized map $r^{\prime}$. Lastly, the argmax function assigns each pixel a cluster label $c_{n}$ based on the maximum value in $r^{\prime}$. Each pixel is assigned the corresponding cluster label $c_{n}$, identical to allocating each pixel to the closest point among the $q^{\prime}$ representative points.
@@
The loss $L$ (defined later) is calculated during backward propagation, and the convolutional filters and classifier parameters are updated using stochastic gradient descent. This process is iterated $T$ times to achieve the final prediction of cluster labels $c_{n}$. The segmentation is handled in an unsupervised manner, with the number of clusters $q^{\prime}$ dynamically updated based on feature similarity and spatial continuity.
@@
3.2 Feature Extractor Network